# Advancing Video Anomaly Detection: A Concise Review and a New Dataset

### v2

@@ Abstract
Video Anomaly Detection (VAD) finds widespread applications in security surveillance, traffic monitoring, industrial monitoring, and healthcare. Despite extensive research efforts, there remains a lack of concise reviews that provide insightful guidance for researchers. Such reviews would serve as quick references to grasp current challenges, research trends, and future directions. In this paper, we present such a review, examining models and datasets from various perspectives. We emphasize the critical relationship between model and dataset, where the quality and diversity of datasets profoundly influence model performance, and dataset development adapts to the evolving needs of emerging approaches. Our review identifies practical issues, including the absence of comprehensive datasets with diverse scenarios.111In this paper, the term ‘scenario’ is used to signify different contexts such as retail, manufacturing, the education sector, smart cities, and others. In the context of different camera views, existing literature sometimes uses the term ‘scene’ to refer to the camera scene from a specific camera viewpoint. To address this, we introduce a new dataset, Multi-Scenario Anomaly Detection (MSAD), comprising 14 distinct scenarios captured from various camera views. Our dataset has diverse motion patterns and challenging variations, such as different lighting and weather conditions, providing a robust foundation for training superior models. We conduct an in-depth analysis of recent representative models using MSAD and highlight its potential in addressing the challenges of detecting anomalies across diverse and evolving surveillance scenarios. Our dataset is available here.
@@ 1 Introduction
Video Anomaly Detection (VAD) aims to automatically identify unusual occurrences in videos, enabling various applications in surveillance and monitoring [82]. Detecting anomalies is a challenging and complex task due to several factors: (i) There is no unified and clear definition of anomalies of interest. For example, the distinction between normal activities like walking on a sidewalk and abnormal activities like walking on a highway is context-dependent. (ii) The sporadic and rare occurrences of anomalies make the collection of well-curated datasets a demanding task, limiting the ability to learn anomalous patterns. Existing methods often treat VAD as either a one-class classification problem or an out-of-distribution detection task [4, 22, 36, 7, 24, 98, 20, 61, 27, 81, 50]. These approaches rely on training solely with normal samples and testing on both normal and abnormal samples, treating anomalies as outliers. Significant progress has been made in VAD during the past few years, thanks to benchmark datasets [35, 79, 37, 72, 14, 97], as well as synthetic data generation [2] that presents a range of anomalies. Fig. 1 shows some frames from different datasets.
@@ 2.1 Dataset deficiencies and biases
Challenges of learned features. Most deep learning-based methods still carefully consider several aspects, such as visual appearance and motion, of human action [33, 20, 27, 86]. While these works highlight the importance of appearance and motion in detecting mostly human-related anomalies [20, 25, 27], non-human-related anomalies remain under-explored due to the lack of solid benchmarks, even when the UCF-Crime dataset (1 non-human-related anomaly) is introduced. Recent works [4] have delved into creating end-to-end deep learning approaches and unified architectures rather than using separate modules or components in a traditional pipeline, as end-to-end solutions are easily accessible, usable, and deployable. However, deep learning solutions require a significant amount of training data, posing a significant concern, especially considering older and smaller datasets such as UCSD Ped [79] and CUHK Avenue [35]. Although efforts have been made in collecting large-scale datasets [72, 92] such as UCF-crime (video- and frame-level annotations for training and testing, respectively), an important issue lies in laborious video annotation, which is one of the main reasons why there haven’t been as many large-scale VAD datasets published yet despite tons of data being publicly available on video sharing sites. A recent emerging few-shot learning framework [19] also encourages researchers to start looking at few-shot VAD models, due to (i) its fast adaptation to novel camera viewpoints/scenarios, (ii) relieving the training data hungry issue, and (iii) its huge potential in real-world applications. Since then, even smaller datasets have proven to be helpful [36, 39].
The beauty of few-shot learning. Few-shot learning aims to adapt quickly to a new task with only a few training samples. One of the most widely recognized methods in VAD is the Few-shot Scene-adaptive Anomaly Detection (FSAD) model [36]. This model uses the meta-learning framework [19] to train a model using video data collected from various camera views within the same scenario, such as ShanghaiTech. The model is then fine-tuned on a different camera viewpoint within the university site (e.g., UCSD Ped and CUHK Avenue). Although the trained model can be adapted to novel viewpoints, its adaptability is still confined to a specific scenario, such as the university street example. The absence of a multi-scenario dataset hampers the widespread application of the rapid adaptation capabilities of VAD models. Recent few-shot VAD methods include [36, 91, 39, 84, 85].
@@
Expanding modalities for human-related anomaly detection. Human-related anomaly detection solutions are not limited to the use of RGB videos. To efficiently extract motions, optical flow, precomputed on RGB videos, has been widely used as temporal or motion information for VAD. Due to the heavy computational cost of optical flow and the redundant nature of RGB videos, researchers have started to explore alternative data modalities for efficient feature extraction [88]. Human pose estimation frameworks, like OpenPose [8], have made human skeleton sequences readily available from RGB videos. This not only introduces a new data modality but also addresses privacy concerns in human-related anomaly detection. The lightweight nature of skeletons and informative spatio-temporal sequences have fostered many skeletal anomaly detection solutions [46, 25]. For example, [46] uses 2D human skeleton motions to detect anomalous human behavior in surveillance footage. Recently, LLMs and pretrained video caption models provide rich descriptions and prompts for video contents to assist VAD tasks [14, 40, 100]. Since then, multi-modal datasets have begun to emerge, with representative examples such as CUVA (RGB videos with text descriptions) in 2024.
Advancements in multi-modal fusion. A growing number of methods have begun integrating multi-modal information in recent years. Relying solely on optical flows [33, 83] or skeletons [24] may not accurately detect anomalies for two main reasons: (i) motions may not reflect relevant anomalies, and (ii) non-human-related anomalies may be treated as background. To address these challenges, several methods have emerged to comprehensively understand anomalies. These include the development of an audio-visual violence dataset and a new fusion method [92], as well as the exploration of semantic information in VAD [12, 53, 93, 99]. One notable approach is a two-branch setup [93] (visual and language-visual branches) that utilizes a pre-trained visual-language model (e.g., CLIP [54]). Additionally, the use of knowledge-based prompts as semantic information [53] and the fusion of text features extracted from a pre-trained video caption model with visual features [12] have been explored. However, compared to single-modality methods, the improvement from these fusion approaches remains somewhat limited, indicating the need for further refinement [62, 12].
@@ 4.1 Generalizability and adaptability
Setup. We first show the superior performance of our scenario-adaptive model by comparing SA2D with the FSAD model [36] on CUHK Avenue. To demonstrate the enhanced anomaly detection performance of our MSAD dataset, we train two adaptive models: one using ShanghaiTech, and the other leveraging our MSAD. For all experiments, we maintain and to ensure a fair comparison. Our training process spans over 1500 epochs. Throughout all evaluations, we consider both Micro and Macro AUC scores. We conduct two sets of experiments: (1) In the single-scenario / cross-view evaluation, the model is trained and tested on the same scenario. To facilitate a more comprehensive comparison with the view-adaptive anomaly detection model [36], we partition the ShanghaiTech dataset into 7 scenes (views), specifically, scene 2, 4, 7, 9, 10, 11, and 12 for training as in [36]. Subsequently, the remaining 5 camera views are individually used for testing purposes.222There are no test videos for scene 13, so we only consider 5 views during testing. (2) In the cross-scenario evaluation, the model is trained on one scenario and tested on a completely different one. For model training, we use either ShanghaiTech or our MSAD.
@@ 4.2 Practical applicability and effectiveness
Setup. We select six recent representative weakly-supervised methods for evaluation: MIST [18], RTFM [76], MSL [31], UR-DMU [102], MGFN [13], and TEVAD [12]. Using Protocol ii, we choose two typical datasets, UCF-Crime and ShanghaiTech, in addition to our MSAD dataset. To maintain consistency and fairness across experiments, we use three popular backbones as video encoders for each method: C3D (2048-dim), I3D (2048-dim, 10-crop), and SwinTransformer (1024-dim). We rigorously maintain original parameters to reproduce reported results from previous studies, and conduct our own evaluations. We use the Micro AUC metric for evaluation.
@@
The role of MSAD. We noticed that our dataset is sensitive to the choice of backbones, particularly for methods like UR-DMU and MGFN. Overall, using I3D as the backbone yields better results compared to using SwinTransformer. RTFM is robust to the choice of backbones, with the performance gap between I3D and SwinTransformer being within 1%. Our MSAD dataset can be used for model backbone selection or for exploring more powerful and practical methods for real-world applications.
@@ 5 Conclusion
Liyun Zhu conducted this research under the supervision of Lei Wang for master’s final year research project at ANU. Liyun Zhu and Arjun Raj are recipients of research sponsorship from Active Intelligence Australia Pty Ltd in Perth, Western Australia, which includes The Active Intelligence Research Challenge Award. This work was also supported by the NCI Adapter Scheme Q4 2023 and the NCI National AI Flagship Merit Allocation Scheme, with computational resources provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.

### v3

@@ Abstract
Video Anomaly Detection (VAD) finds widespread applications in security surveillance, traffic monitoring, industrial monitoring, and healthcare. Despite extensive research efforts, there remains a lack of concise reviews that provide insightful guidance for researchers. Such reviews would serve as quick references to grasp current challenges, research trends, and future directions. In this paper, we present such a review, examining models and datasets from various perspectives. We emphasize the critical relationship between model and dataset, where the quality and diversity of datasets profoundly influence model performance, and dataset development adapts to the evolving needs of emerging approaches. Our review identifies practical issues, including the absence of comprehensive datasets with diverse scenarios.111In this paper, the term ‘scenario’ is used to signify different contexts such as retail, manufacturing, the education sector, smart cities, and others. In the context of different camera views, existing literature sometimes uses the term ‘scene’ to refer to the camera scene from a specific camera viewpoint. To address this, we introduce a new dataset, Multi-Scenario Anomaly Detection (MSAD), comprising 14 distinct scenarios captured from various camera views. Our dataset has diverse motion patterns and challenging variations, such as different lighting and weather conditions, providing a robust foundation for training superior models. We conduct an in-depth analysis of recent representative models using MSAD and highlight its potential in addressing the challenges of detecting anomalies across diverse and evolving surveillance scenarios. [Project website]
@@ 1 Introduction
Video Anomaly Detection (VAD) aims to automatically identify unusual occurrences in videos, enabling various applications in surveillance and monitoring [92]. Detecting anomalies is a challenging and complex task due to several factors: (i) There is no unified and clear definition of anomalies of interest.222Although there is no unified and clear definition of specific anomalies. Generally, we improve the definition of anomaly in the context of video anomaly detection, an anomaly refers to any event, behavior, or object in a video sequence that deviates from the normal or expected pattern of events. Given that anomalies are highly scene-dependent, single-scene datasets have a distinct advantage in testing a model’s performance under consistent, controlled conditions. This allows for a more focused evaluation of the model’s ability to detect deviations within a specific context. For example, the distinction between normal activities like walking on a sidewalk and abnormal activities like walking on a highway is context-dependent. (ii) The sporadic and rare occurrences of anomalies make the collection of well-curated datasets a demanding task, limiting the ability to learn anomalous patterns. Existing methods often treat VAD as either a one-class classification problem or an out-of-distribution detection task [4, 25, 42, 7, 28, 113, 23, 68, 31, 88, 56]. These approaches rely on training solely with normal samples and testing on both normal and abnormal samples, treating anomalies as outliers. Significant progress has been made in VAD during the past few years, thanks to benchmark datasets [41, 86, 43, 79, 17, 112], as well as synthetic data generation [2] that presents a range of anomalies. Fig. 1 shows some frames from different datasets.
@@ 2.1 Dataset deficiencies and biases
Challenges of learned features. Most deep learning-based methods still carefully consider several aspects, such as visual appearance and motion, of human action [89, 39, 31, 98, 23, 93, 101, 91, 90, 12]. While these works highlight the importance of appearance and motion in detecting mostly human-related anomalies [23, 29, 31], non-human-related anomalies remain under-explored due to the lack of solid benchmarks, even when the UCF-Crime dataset (1 non-human-related anomaly) is introduced. Recent works [4] have delved into creating end-to-end deep learning approaches and unified architectures rather than using separate modules or components in a traditional pipeline, as end-to-end solutions are easily accessible, usable, and deployable. However, deep learning solutions require a significant amount of training data, posing a significant concern, especially considering older and smaller datasets such as UCSD Ped [86] and CUHK Avenue [41]. Although efforts have been made in collecting large-scale datasets [79, 107] such as UCF-crime (video- and frame-level annotations for training and testing, respectively), an important issue lies in laborious video annotation, which is one of the main reasons why there haven’t been as many large-scale VAD datasets published yet despite tons of data being publicly available on video sharing sites. A recent emerging few-shot learning framework [22] also encourages researchers to start looking at few-shot VAD models, due to (i) its fast adaptation to novel camera viewpoints/scenarios, (ii) relieving the training data hungry issue, and (iii) its huge potential in real-world applications. Since then, even smaller datasets have proven to be helpful [42, 45].
The beauty of few-shot learning. Few-shot learning aims to adapt quickly to a new task with only a few training samples [99, 94, 95, 100]. One of the most widely recognized methods in VAD is the Few-shot Scene-adaptive Anomaly Detection (FSAD) model [42]. This model uses the meta-learning framework [22] to train a model using video data collected from various camera views within the same scenario, such as ShanghaiTech. The model is then fine-tuned on a different camera viewpoint within the university site (e.g., UCSD Ped and CUHK Avenue). Although the trained model can be adapted to novel viewpoints, its adaptability is still confined to a specific scenario, such as the university street example. The absence of a multi-scenario dataset hampers the widespread application of the rapid adaptation capabilities of VAD models. Recent few-shot VAD methods include [42, 106, 45, 99, 100].
@@
Expanding modalities for human-related anomaly detection. Human-related anomaly detection solutions are not limited to the use of RGB videos. To efficiently extract motions, optical flow, precomputed on RGB videos, has been widely used as temporal or motion information for VAD. Due to the heavy computational cost of optical flow and the redundant nature of RGB videos, researchers have started to explore alternative data modalities for efficient feature extraction [103]. Human pose estimation frameworks, like OpenPose [8], have made human skeleton sequences readily available from RGB videos. This not only introduces a new data modality but also addresses privacy concerns in human-related anomaly detection. The lightweight nature of skeletons [60, 34, 96] and informative spatio-temporal sequences have fostered many skeletal anomaly detection solutions [52, 29]. For example, [52] uses 2D human skeleton motions to detect anomalous human behavior in surveillance footage. Recently, LLMs and pretrained video caption models provide rich descriptions and prompts for video contents to assist VAD tasks [17, 46, 115]. Since then, multi-modal datasets have begun to emerge, with representative examples such as CUVA (RGB videos with text descriptions) in 2024.
Advancements in multi-modal fusion. A growing number of methods have begun integrating multi-modal information in recent years. Relying solely on optical flows [39, 97] or skeletons [28] may not accurately detect anomalies for two main reasons: (i) motions may not reflect relevant anomalies, and (ii) non-human-related anomalies may be treated as background. To address these challenges, several methods have emerged to comprehensively understand anomalies. These include the development of an audio-visual violence dataset and a new fusion method [107, 15], as well as the exploration of semantic information in VAD [13, 59, 108, 114]. One notable approach is a two-branch setup [108] (visual and language-visual branches) that utilizes a pre-trained visual-language model (e.g., CLIP [61]). Additionally, the use of knowledge-based prompts as semantic information [59] and the fusion of text features extracted from a pre-trained video caption model with visual features [13] have been explored. However, compared to single-modality methods, the improvement from these fusion approaches remains somewhat limited, indicating the need for further refinement [69, 13].
@@ 4.1 Generalizability and adaptability
Setup. We first show the superior performance of our scenario-adaptive model by comparing SA2D with the FSAD model [42] on CUHK Avenue. To demonstrate the enhanced anomaly detection performance of our MSAD dataset, we train two adaptive models: one using ShanghaiTech, and the other leveraging our MSAD. For all experiments, we maintain and to ensure a fair comparison. Our training process spans over 1500 epochs. Throughout all evaluations, we consider both Micro and Macro AUC scores. We conduct two sets of experiments: (1) In the single-scenario / cross-view evaluation, the model is trained and tested on the same scenario. To facilitate a more comprehensive comparison with the view-adaptive anomaly detection model [42], we partition the ShanghaiTech dataset into 7 scenes (views), specifically, scene 2, 4, 7, 9, 10, 11, and 12 for training as in [42]. Subsequently, the remaining 5 camera views are individually used for testing purposes.333There are no test videos for scene 13, so we only consider 5 views during testing. (2) In the cross-scenario evaluation, the model is trained on one scenario and tested on a completely different one. For model training, we use either ShanghaiTech or our MSAD.
@@
Scene information is implicitly used as weak supervision. The scene information provided is used in the few-shot sampling strategy, where each scenario is treated as a group for sampling purposes. This allows the model to learn information from multiple scenarios in a balanced manner (see Fig. 6 in Appendix D). In our proposed model, SA2D, scene information is used during the sampling process to form -way -shot learning. Table 3’s last four columns compare the performance (i) without scene information (FSAD) and (ii) with scene information (our SA2D) on our MSAD dataset (note that both models are trained using MSAD). The primary difference between FSAD and SA2D is that SA2D uses scene information during sampling to enhance the -way -shot learning framework. As shown in Table 3, our sampling strategy (Appendix D) significantly boosts SA2D’s performance across all five test splits on different camera viewpoints in ShanghaiTech.
@@ 4.2 Practical applicability and effectiveness
Setup. We focus on weakly-supervised learning methods from a practical perspective. We select six recent representative methods for evaluation: MIST [21], RTFM [83], MSL [37], UR-DMU [117], MGFN [14], and TEVAD [13]. These methods provide open-source implementations, allowing us to explore the impact of different backbone choices. Each method has its unique focus. MIST introduces a multiple instance self-training framework to refine task-specific discriminative representations using only video-level annotations. RTFM trains a feature magnitude learning function to identify positive instances, enhancing the robustness of the multiple instance learning approach. MSL uses multi-sequence learning to predict both video-level anomaly probabilities and snippet-level anomaly scores. UR-DMU focuses on learning representations of normal data while extracting discriminative features from abnormal data for a better understanding of normal states. MGFN proposes a glance-and-focus network to amplify the discriminative power of feature magnitudes across different scenes. TEVAD uses both visual and text features to complement spatio-temporal features with semantic meanings of abnormal events. These methods enable us to conduct a fair and comprehensive evaluation across various aspects.
In line with the standard practice in [83, 37, 117], we use C3D [84], I3D [9], SwinTransformer [40] pretrained on Kinetics-400, as backbone networks for feature extraction. Although the same backbone is used, different methods may extract features of varying dimensions. For example, UR-DMU and MSL extracts 1024-dim features, whereas other methods extract 2048-dim features when using I3D as the backbone. To ensure a fair comparison, following RTFM, we extract 4096-dim features from the ‘fc6’ layer of C3D, 2048-dim, 10-crop features from the ‘mix 5c’ layer of I3D, respectively. Additionally, following MSL, we extract 1024-dim features from SwinTransformer. We apply these feature extraction processes on ShanghaiTech, UCF-Crime and our MSAD. Then we use these extracted features as inputs of different methods. We rigorously maintain original parameters to reproduce reported results from previous studies, and conduct our own evaluations using Protocol ii. We use the Micro AUC metric for evaluation.
@@
MSAD is sensitive to the choice of backbones. We notice that our dataset is sensitive to the choice of backbones, particularly for methods like UR-DMU and MGFN (Table 4). Existing backbones, such as C3D, I3D, and SwinTransformer, have demonstrated strong performance in video anomaly detection on current datasets (see Table 4 for results on UCF-Crime and ShanghaiTech). However, these datasets have significant limitations: (i) they cover a very limited range of scenarios, motion dynamics, and anomaly types, and (ii) most anomaly detection methods rely on action recognition pre-trained models as backbones for feature extraction, particularly for human-related anomalies. As a result, these backbones tend to perform well because the anomaly detection datasets predominantly feature human-related motions. In contrast, our dataset introduces more challenges in terms of scenarios, camera viewpoints, and motion dynamics, including both human and non-human-related motions. Overall, using I3D as the backbone yields better results compared to using SwinTransformer. RTFM is robust to the choice of backbones, with the performance gap between I3D and SwinTransformer being within 1%.
Our dataset can be used for selecting appropriate model backbones or exploring more powerful backbone networks that do not overly depend on existing action recognition models. Additionally, our dataset advances video anomaly detection by considering a wider range of scenarios and a broader spectrum of anomaly types.
4.3 Limitations of current methods on our new dataset
Here, we discuss the limitations of current methods and provide additional insights on our MSAD dataset.
I3D vs. SwinTransformer. Using I3D as a backbone generally outperforms SwinTransformer. Although SwinTransformer excels in capturing local and global spatial features, its reliance on attention mechanisms for temporal modeling may not be as finely tuned to detect subtle anomalies. I3D, designed to maintain temporal consistency across frames, is better suited for detecting anomalies that are temporally localized, e.g., sudden object appearances or unexpected behaviors. SwinTransformer’s attention-based approach may lack the temporal coherence needed for such tasks. Additionally, I3D’s focus on motion and spatiotemporal features aligns well with anomaly detection requirements, especially in motion-based anomalies and subtle temporal variations present in our dataset.
Boosting existing methods. Models trained on our dataset exhibit better generalizability and adaptability, particularly under few-shot settings (Table 3), showing superior performance on unseen camera viewpoints. Our SA2D model, trained on MSAD (Table 3), significantly boosts performance. Additionally, our dataset enhances the performance of existing anomaly detection methods (Table 7 in the Appendix). Models trained on MSAD consistently achieve the best performance on anomalies such as fighting, fire, object falling, shooting, traffic accidents, and water incidents, addressing the challenge of lacking a comprehensive benchmark dataset for model training.
No single best performer on MSAD. RTFM and TEVAD are more robust to backbone changes than UR-DMU and MGFN (see Table 4). This suggests that existing methods may lack comprehensive evaluations across diverse scenarios. Our dataset provides a solid benchmark for training, testing, and evaluating superior models.
Supporting multi-scenario anomaly detection. Our dataset benefits multi-scenario anomaly detection, enabling systematic evaluation across multiple scenarios (Table 8 in the Appendix). Recent methods trained on MSAD generally achieve better performance across all 14 scenarios, demonstrating that existing works have primarily focused on single-scenario detection. Our dataset offers a robust foundation for training superior models for multi-scenario anomaly evaluation.
Rethinking anomaly detection. Our dataset encourages rethinking the correctness and trustworthiness of anomaly detection methods. Cross-dataset evaluation (Table 9 in the Appendix) reveals that existing datasets (Table 1) may suffer from unrealistic anomaly definitions. For instance, models pre-trained on MSAD perform poorly on ShanghaiTech but better on CUHK and UCSD Ped2, highlighting the misalignment of anomaly categories like biking and driving with reality.
@@ 5 Conclusion
Liyun Zhu conducted this research under the supervision of Lei Wang for his final year master’s research project at ANU. Liyun Zhu and Arjun Raj are recipients of The Active Intelligence Research Challenge Award. This work was also supported by the NCI Adapter Scheme Q4 2023 and the NCI National AI Flagship Merit Allocation Scheme, with computational resources provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
Checklist
The checklist follows the references. Please read the checklist guidelines carefully for information on how to answer these questions. For each question, change the default [TODO] to [Yes] , [No] , or [N/A] . You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description. For example:
Did you include the license to the code and datasets? [Yes]
Did you include the license to the code and datasets? [No] The code and the data are proprietary.
Did you include the license to the code and datasets? [N/A]
Please do not modify the questions and only use the provided macros for your answers. Note that the Checklist section does not count towards the page limit. In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
i.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes]
Did you discuss any potential negative societal impacts of your work? [Yes]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
ii.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
iii.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
iv.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [Yes]
Did you include any new assets either in the supplemental material or as a URL? [Yes]
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes]
v.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]